While standing at the edge of the roof of a building, a man throws a stone upward with an initial speed of 7.25 m/s. The stone subsequently falls to the ground, which is 12.1 m below the point where the stone leaves his hand. At what speed does the stone impact the ground? Ignore air resistance and use g 9.81 m/s2 for the acceleration due to gravity. impact speed: m/s How much time is the stone in the air? elapsed time:
Answer #1

a)

using 3rd equation of motion

v^2 = u^2 + 2 g h

v^2 = 7.25^2 + 2* 9.81* 12.1

v = 17.03 m/s

=========

b)

v = - u + gt

17.03 = - 7.25 + 9.81* t

t = 2.475 seconds

========

do comment in case any doubt, will reply for sure.. Goodluck